# Autonomous Navigation and Station-Keeping on Near-Rectilinear Halo Orbits

### v1

@@ 3 Autonomous Navigation with Optical Measurements
We first provide an overview of the horizon-based OPNAV scheme. We adopt the non-iterative Christian-Robinson algorithm [4] due to its superior performance over other iterative, ellipse-fitting-based variants [25]. We develop a measurement covariance expression based on [4] which accounts for attitude uncertainty in addition to the error associated with detecting the lit limb within the image. Then, we briefly discuss the simulation environment developed in Blender in our previous work [21]. Finally, we describe the dynamic filter employed with the OPNAV measurements for simulating autonomous navigation.
@@ 3.1 Horizon-Based Optical Navigation
We consider the horizon-based OPNAV algorithm for computing the position vector of the spacecraft with respect to the imaged body assuming the camera attitude, defined by the camera frame $\mathcal{F}_{\rm C}$, is known. This assumption is built on the fact that spacecraft attitude can be computed with high precision using star trackers. The small error on the spacecraft’s attitude is incorporated into the image generation process, and the measurement covariance expression is updated from the expression provided in [4] to reflect this source of error. The algorithm also requires ephemerides of the Moon and the Sun, as well as the radii of the Moon; these are assumed to be stored on the onboard computer.
@@ 3.1.1 Christian-Robinson Algorithm
The Christian-Robinson algorithm [4] provides both the position vector estimate as well as an analytical expression for the position measurement covariance. While a short overview is provided here for completeness, a full description is available for the Cholesky decomposition variant in [4] and for the Singular Value Decomposition (SVD) variant in [5]. Starting with the $m$ pixel coordinates along the lit limb $\{u_{i},v_{i}\}_{i=1}^{m}$ of the Moon, each point is projected onto a point on the image plane $\boldsymbol{s}_{i}$ by using the intrinsic camera calibration matrix $\boldsymbol{\kappa}$ [26, 5, 6]
$\boldsymbol{s}_{i}=\boldsymbol{\kappa}^{-1}\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix}$. (5)
Then, using the SVD, the horizon-based OPNAV problem is transformed from being relative to a triaxial ellipsoid to being relative to a unit sphere via
@@
$\begin{bmatrix}{\bar{\boldsymbol{s}}^{\prime T}_{1}}\\ {\bar{\boldsymbol{s}}^{\prime T}_{2}}\\ \vdots\\ {\bar{\boldsymbol{s}}^{\prime T}_{m}}\end{bmatrix}\boldsymbol{n}=\boldsymbol{H}\boldsymbol{n}=\boldsymbol{1}_{m\times 1}.$ (7)
The matrix $\boldsymbol{H}$ is the vertical concatenation of $\bar{\boldsymbol{s}}^{\prime T}_{i}$} for $i=1,\ldots,m$. The position vector of the spacecraft in $\mathcal{F}_{\rm C}$ is given by
@@ 3.1.2 Analytical Measurement Covariance Under Attitude Uncertainty
Following Christian and Robinson [4], we compute the covariance of the position measurement, in the presence of error associated with detecting the lit limb within the image, quantified in terms of the “standard deviation of an observed horizon point in units of pixels” $\sigma_{\mathrm{pix}}$. This model assumes a perfect attitude knowledge, and as a consequence a perfect knowledge of the transformation matrix $\boldsymbol{T}^{\rm P}_{\rm C}$. However, in reality, the camera will have some unknown misalignment due to the imperfection of the estimated attitude, the mounting process of the hardware, and the vibration during launch. To this end, we develop an expression for the measurement covariance of the position in $\mathcal{F}_{\rm P}$, denoted by $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{P}}}$, that incorporates both the error introduced by the limb detection process, as well as the imperfect attitude knowledge.
Taking the variation of (9) with respect to both $\boldsymbol{n}$ and $\boldsymbol{\phi}$,
@@
An approximation for $\boldsymbol{P}_{n}$ developed in [27] for Cholesky decomposition is adapted to the SVD via
@@
$\sigma_{\mathrm{pix}}$ is the standard deviation of an observed horizon point in units of pixels [27], and $\boldsymbol{J}_{i}$ is given by
@@
    $\boldsymbol{r}_{\mathrm{P}}=\delta\boldsymbol{T}\boldsymbol{T}^{\rm C}_{\rm P}\begin{bmatrix}0\\ 0\\ -70000\end{bmatrix}$. (21)
@@
    Compute position estimate $\hat{\boldsymbol{r}}_{\mathrm{P}}$ and covariance $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{P}}}$ from Christian-Robinson algorithm using $\boldsymbol{T}^{\rm C}_{\rm P}$.
@@
    Compute error as $\hat{\boldsymbol{r}}_{\mathrm{P}}-\boldsymbol{r}_{\mathrm{P}}$.
Using $\sigma_{\rm pix}=0.5$ pixel and $\sigma_{\phi}=15$ $\mathrm{arcsec}$, the distribution of the actual measurement error is shown in Figure 2. Based on the $10\,000\text{\,}$ samples, we find $66.060\%$ fall within $1$-$\sigma$, $94.710\%$ fall within $2$-$\sigma$, and $99.730\%$ fall within $3$-$\sigma$ as predicted by $\boldsymbol{P}_{\boldsymbol{r}_{\rm P}}$.
Figure 2: Monte-Carlo results for measurement distributions of horizon-based OPNAV position under attitude error, using 10000 tries with $\sigma_{\rm pix}=0.5$ pixel, $\sigma_{\phi}=15$ $\mathrm{arcsec}$. Red lines indicate empirical 22-$\sigma$ of the samples.
@@ 3.2 Simulation Environment for Horizon-Based Optical Navigation
The simulation environment is developed using Blender [28], an open-source rendering engine. The development process of the simulation environment is detailed in our prior paper [21]. Points on the image corresponding to the lit limb of the Moon are found through a three-step process following [5], consisting of (i) creating a mask centered around approximate lit limb locations found by scanning the image along the projected direction of the illumination vector of the Sun, followed by (ii) applying a mask on edge points detected using the Canny edge detection algorithm, and finally (iii) applying a Zernike moment-based refinement scheme to correct the pixel-level point from (ii) down to subpixel-level locations using local mini patches around each point.
@@ 3.3 Navigation Filter
In the context of spaceflight, the EKF has been successfully applied in multiple missions [29]. Through preliminary studies, we found that EKF and UKF result in nearly identical navigation performance assuming a measurement update cadence of a few images per revolution. Thus, we adopt the EKF for navigation. Nevertheless, we also introduce the UT-based prediction, as its prediction scheme is leveraged for the design of the SK controller.
@@ 3.3.1 Prediction in Extended Kalman Filter
To account for the uncertainties in (1), we include an additive process noise [30]. The prediction step of the EKF is given by
$\hat{\boldsymbol{\theta}}_{j\mid j-1}$ $=\hat{\boldsymbol{\theta}}_{j-1\mid j-1}+\int_{t_{j-1}}^{t_{j}}\boldsymbol{f}(\tau,\hat{\boldsymbol{\theta}}_{j-1}(\tau))\mathrm{d}\tau,$ (22a)
@@
The process noise is modeled as an unbiased random process with covariance
@@
where $h_{j}=t_{j}-t_{j-1}$ is the duration of a sampling period and $\sigma_{u}^{2}$ is the diffusion coefficient to be adjusted to optimize the filter performance [31].
@@ 3.3.2 Prediction in Unscented Kalman Filter
In the UKF, the prediction involves forming and propagating $2n+1$ sigma points, chosen as in [32],
$\mathcal{X}_{j-1\mid j-1}^{(0)}$ $=\hat{\boldsymbol{\theta}}_{j-1\mid j-1},$ (24a)
$\mathcal{X}_{j-1\mid j-1}^{(\ell)}$ $=\hat{\boldsymbol{\theta}}_{j-1\mid j-1}+\sqrt{n+\lambda}\left[\sqrt{\boldsymbol{\Sigma}_{j-1\mid j-1}}\right]_{\ell},\quad\ell=1,\ldots,n,$ (24b)
$\mathcal{X}_{j-1\mid j-1}^{(\ell+n)}$ $=\hat{\boldsymbol{\theta}}_{j-1\mid j-1}-\sqrt{n+\lambda}\left[\sqrt{\boldsymbol{\Sigma}_{j-1\mid j-1}}\right]_{\ell},\quad\ell=1,\ldots,n,$ (24c)
@@
$\mathcal{X}_{j\mid j-1}^{(i)}=\mathcal{X}_{j-1\mid j-1}^{(\ell)}+\int_{t_{j-1}}^{t_{j}}f(\tau,\boldsymbol{\theta}(\tau))\mathrm{d}\tau,\quad\ell=0,\ldots,n.$ (26)
@@
$\hat{\boldsymbol{\theta}}_{j\mid j-1}$ $=\sum_{\ell=0}^{2n}W_{\ell}^{(m)}\mathcal{X}_{j\mid j-1}^{(\ell)},$ (27a)
$\boldsymbol{\Sigma}_{j\mid j-1}$ $=\sum_{\ell=0}^{2n}W_{\ell}^{(c)}\left(\mathcal{X}_{j\mid j-1}^{(\ell)}-\hat{\boldsymbol{\theta}}_{j\mid j-1}\right)\left(\mathcal{X}_{j\mid j-1}^{(\ell)}-\hat{\boldsymbol{\theta}}_{j\mid j-1}\right)^{T}+\boldsymbol{Q}_{j},$ (27b)
where $\boldsymbol{Q}_{j}$ is the process noise given by expression (23), while $W_{i}^{(m)}$ and $W_{\ell}^{(c)}$ are weights given by
$W_{0}^{(m)}$ $=\dfrac{\lambda}{n+\lambda},$ (28)
$W_{0}^{(c)}$ $=\dfrac{\lambda}{n+\lambda+(1-\alpha^{2}+\beta)},$
$W_{\ell}^{(m)}$ $=\dfrac{1}{2(n+\lambda)},\quad\ell=1,\ldots,2n,$
$W_{\ell}^{(c}$ $=\dfrac{1}{2(n+\lambda)},\quad\ell=1,\ldots,2n,$
and $\beta$ is an additional parameter to incorporate prior knowledge of $\hat{\boldsymbol{\theta}}$; $\beta=2$ is optimal assuming the distribution of $\hat{\boldsymbol{\theta}}$ is Gaussian.

### v2

@@ 3 Autonomous Navigation with Optical Measurements
We first provide an overview of the horizon-based OPNAV scheme. We adopt the non-iterative Christian-Robinson algorithm [ChristianRobinson2016] due to its superior performance over other iterative, ellipse-fitting-based variants [Christian2016Iterative]. We develop a measurement covariance expression based on [ChristianRobinson2016] which accounts for attitude uncertainty in addition to the error associated with detecting the lit limb within the image. Then, we briefly discuss the simulation environment developed in Blender in our previous work [Shimane2023oct]. Finally, we describe the dynamic filter employed with the OPNAV measurements for simulating autonomous navigation.
@@ 3.1 Horizon-Based Optical Navigation
We consider the horizon-based OPNAV algorithm for computing the position vector of the spacecraft with respect to the imaged body assuming the camera attitude, defined by the camera frame $\mathcal{F}_{\rm C}$, is known. This assumption is built on the fact that spacecraft attitude can be computed with high precision using star trackers. The small error on the spacecraft’s attitude is incorporated into the image generation process, and the measurement covariance expression is updated from the expression provided in [ChristianRobinson2016] to reflect this source of error. The algorithm also requires ephemerides of the Moon and the Sun, as well as the radii of the Moon; these are assumed to be stored on the onboard computer.
@@ 3.1.1 Christian-Robinson Algorithm
The Christian-Robinson algorithm [ChristianRobinson2016] provides both the position vector estimate as well as an analytical expression for the position measurement covariance in the body’s principal axes frame $\mathcal{F}_{\rm P}$. While a short overview is provided here for completeness, a full description is available for the Cholesky decomposition variant in [ChristianRobinson2016] and for the Singular Value Decomposition (SVD) variant in [Christian2017]. Starting with the $m$ pixel coordinates along the lit limb $\{u_{i},v_{i}\}_{i=1}^{m}$ of the Moon and assuming a pinhole camera model [Christian2021Tutorial], each point is projected onto a point on the image plane $\boldsymbol{s}_{i}$ by using the intrinsic camera calibration matrix $\boldsymbol{\kappa}$ [Hartley2003, Christian2017, Christian2021Tutorial]
$\boldsymbol{s}_{i}=\boldsymbol{\kappa}^{-1}\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix},\quad\boldsymbol{\kappa}=\begin{bmatrix}f/\mu_{x}&\alpha&u_{p}\\ 0&f/\mu_{y}&v_{p}\\ 0&0&1\end{bmatrix}$, (7)
where $f$ is the focal length, $\mu_{x}$ and $\mu_{y}$ are the horizontal and vertical pixel pitch, $u_{p},v_{p}]$ defines the principal point of the focal plane, and $\alpha$ is the skewness. In this work, we assume the synthetic images contain no affine shear, i.e., $\alpha=0$. Here, we construct $\boldsymbol{\kappa}$ using the exact values for $f$, $\mu_{x}$, $\mu_{y}$, $u_{p}$, and $v_{p}$ used to generate the synthetic images of the Moon. While in practice, errors in $\boldsymbol{\kappa}$ are inevitable, their impact is mitigated through on-orbit re-calibration of $\boldsymbol{\kappa}$ at the early stage of a mission [Christian2016-qp]. In this work, we choose to assume a perfect $\boldsymbol{\kappa}$, thereby isolating the performance of the station-keeping pipeline from errors introduced by the optical hardware. If needed, these could be accounted for by appropriately increasing the measurement error covariance. Using SVD, the horizon-based OPNAV problem is transformed from being relative to a triaxial ellipsoid to being relative to a unit sphere via
@@
$\begin{bmatrix}{\bar{\boldsymbol{s}}^{\prime T}_{1}}\\ {\bar{\boldsymbol{s}}^{\prime T}_{2}}\\ \vdots\\ {\bar{\boldsymbol{s}}^{\prime T}_{m}}\end{bmatrix}\boldsymbol{n}=\boldsymbol{H}\boldsymbol{n}=\boldsymbol{1}_{m\times 1},$ (9)
where $\boldsymbol{n}$ is a vector in the direction of $\boldsymbol{r}$ with a length related to the distance between the camera and the center of the imaged body [Christian2021_TRN]. The matrix $\boldsymbol{H}$ is the vertical concatenation of $\bar{\boldsymbol{s}}^{\prime T}_{i}$} for $i=1,\ldots,m$. The position vector of the spacecraft in $\mathcal{F}_{\rm C}$ is given by
@@ 3.1.2 Analytical Measurement Covariance Under Attitude Uncertainty
Following Christian and Robinson [ChristianRobinson2016], we compute the covariance of the position measurement, in the presence of error associated with detecting the lit limb within the image, quantified in terms of the “standard deviation of an observed horizon point in units of pixels” $\sigma_{\mathrm{pix}}$. This model assumes a perfect attitude knowledge, and as a consequence a perfect knowledge of the transformation matrix $\boldsymbol{T}^{\rm P}_{\rm C}$. However, in reality, the camera will have some unknown misalignment due to the imperfection of the estimated attitude, the mounting process of the hardware, and the vibration during launch. To this end, we develop an expression for the measurement covariance of the position in $\mathcal{F}_{\rm P}$, denoted by $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{P}}}$, that incorporates both the error introduced by the limb detection process, as well as the imperfect attitude knowledge.
Let $\boldsymbol{\phi}$ denote the rotation vector corresponding to $\boldsymbol{T}^{\rm C}_{\rm P}$. Taking the variation of (11) with respect to both $\boldsymbol{n}$ and $\boldsymbol{\phi}$,
@@
An approximation for $\boldsymbol{P}_{n}$ developed in [Christian2016] for Cholesky decomposition is adapted to the SVD via
@@
$\sigma_{\mathrm{pix}}$ is the standard deviation of an observed horizon point in units of pixels [Christian2016], and $\boldsymbol{J}_{i}$ is given by
@@
    $\boldsymbol{r}_{\mathrm{P}}=\delta\boldsymbol{T}\boldsymbol{r}_{\rm C},\quad\boldsymbol{r}_{\rm C}=\boldsymbol{T}^{\rm C}_{\rm P}\begin{bmatrix}0\\ 0\\ -70000\end{bmatrix}$. (23)
@@
    Compute position estimate $\hat{\boldsymbol{r}}_{\mathrm{C}}$ and covariance $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{C}}}$ with Christian-Robinson algorithm and convert to $\mathcal{F}_{\rm P}$, $\hat{\boldsymbol{r}}_{\mathrm{P}}$ and $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{P}}}$, using transformation matrix $\boldsymbol{T}^{\rm C}_{\rm P}$.
@@
    Compute error in $\mathcal{F}_{\rm C}$ as $\hat{\boldsymbol{r}}_{\mathrm{C}}-\boldsymbol{r}_{\mathrm{C}}$ and in $\mathcal{F}_{\rm P}$ as $\hat{\boldsymbol{r}}_{\mathrm{P}}-\boldsymbol{r}_{\mathrm{P}}$.
Using $\sigma_{\rm pix}=0.5$ pixel and $\sigma_{\phi}=15$ $\mathrm{arcsec}$, the distribution of the actual measurement error based on 10000 samples is shown in Figure 2 in $\mathcal{F}_{\rm C}$ and in $\mathcal{F}_{\rm P}$. In the Figure 2(a), the red lines indicate $2$-$\sigma$ calculated from $\boldsymbol{P}_{\boldsymbol{r}_{\rm C}}$ of the samples. Based on the 10000 samples, we find $71.350\%$ fall within $1$-$\sigma$, $95.740\%$ fall within $2$-$\sigma$, and $99.670\%$ fall within $3$-$\sigma$ as predicted by $\boldsymbol{P}_{\boldsymbol{r}_{\rm P}}$.
(a) Error in Camera frame
(b) Error in randomized Moon principal axes frame
Figure 2: Measurement error distributions of horizon-based OPNAV position under attitude error
@@ 3.2 Simulation Environment for Horizon-Based Optical Navigation
The simulation environment is developed using Blender [Blender], an open-source rendering engine. The development process of the simulation environment is detailed in our prior paper [Shimane2023oct]. Points on the image corresponding to the lit limb of the Moon are found through a three-step process following [Christian2017], consisting of (i) creating a mask centered around approximate lit limb locations found by scanning the image along the projected direction of the illumination vector of the Sun, followed by (ii) applying a mask on edge points detected using the Canny edge detection algorithm, and finally (iii) applying a Zernike moment-based refinement scheme to correct the pixel-level point from (ii) down to subpixel-level locations using local mini patches centered around each pixel-level edge. Figure 3 shows, from left to right, the original synthetic image of the Moon, the illumination scan projected into the image, the masks created by the illumination scan, and the pixel-level and subpixel-level refined edges.
Figure 3: Synthetic image, illumination scan, masked regions, and subpixel-level edge detection
@@ 3.3 Navigation Filter
In the context of spaceflight, the EKF has been successfully applied in multiple missions [Carpenter2018]. Through preliminary studies, we found that EKF and UKF result in nearly identical navigation performance assuming a measurement update cadence of a few images per revolution. Thus, we adopt the EKF for navigation. Nevertheless, we also introduce the UT-based prediction, as its prediction scheme is leveraged for the design of the SK controller.
@@ 3.3.1 Prediction in Extended Kalman Filter
The prediction step consists of propagating (1), accounting for uncertainties in the SRP term (3). The filter assumes nominal SRP coefficients $\bar{C}_{r}$ and $\overline{A/m}$, while their true values are perturbed as
$C_{r}$ $=\bar{C}_{r}+(1+\delta_{C_{r}}),\quad\delta_{C_{r}}\sim\mathcal{N}(0,\sigma_{C_{r}}^{2}),\quad$ (24a)
$A/m$ $=\overline{A/m}(1+\delta_{A/m}),\quad\delta_{A/m}\sim\mathcal{N}(0,\sigma_{A/m}^{2}),$ (24b)
at the beginning of each prediction. The uncertainty is accounted for by an additive process noise [sarkka2023bayesian], The prediction step of the EKF is given by
$\hat{\boldsymbol{x}}_{j\mid j-1}$ $=\hat{\boldsymbol{x}}_{j-1\mid j-1}+\int_{t_{j-1}}^{t_{j}}\boldsymbol{f}(\tau,\hat{\boldsymbol{x}}_{j-1}(\tau))\mathrm{d}\tau,$ (25a)
@@
$\boldsymbol{Q}_{j}$ is the additive process noise that accounts for the dynamics modeling error [sarkka2023bayesian], modeled as an unbiased random process
@@
where $h_{j}=t_{j}-t_{j-1}$ is the duration of a sampling period and $\sigma_{u}^{2}$ is the diffusion coefficient to be adjusted to optimize the filter performance [Tapley_Schutz_Born_2004].
@@ 3.3.2 Prediction in Unscented Kalman Filter
In the UKF, the prediction involves forming and propagating $2n+1$ sigma points, chosen as in [Sarkka2007],
$\mathcal{X}_{j-1\mid j-1}^{(0)}$ $=\hat{\boldsymbol{x}}_{j-1\mid j-1},$ (29a)
$\mathcal{X}_{j-1\mid j-1}^{(\ell)}$ $=\hat{\boldsymbol{x}}_{j-1\mid j-1}+\sqrt{n+\lambda}\left[\sqrt{\boldsymbol{\Sigma}_{j-1\mid j-1}}\right]_{\ell},\quad\ell=1,\ldots,n,$ (29b)
$\mathcal{X}_{j-1\mid j-1}^{(\ell+n)}$ $=\hat{\boldsymbol{x}}_{j-1\mid j-1}-\sqrt{n+\lambda}\left[\sqrt{\boldsymbol{\Sigma}_{j-1\mid j-1}}\right]_{\ell},\quad\ell=1,\ldots,n,$ (29c)
@@
$\mathcal{X}_{j\mid j-1}^{(\ell)}=\mathcal{X}_{j-1\mid j-1}^{(\ell)}+\int_{t_{j-1}}^{t_{j}}f\left(\tau,\mathcal{X}_{j-1\mid j-1}^{(\ell)}(\tau)\right)\mathrm{d}\tau,\quad\ell=0,\ldots,n.$ (31)
@@
$\hat{\boldsymbol{x}}_{j\mid j-1}$ $=\sum_{\ell=0}^{2n}W^{(\ell)}_{m}\mathcal{X}_{j\mid j-1}^{(\ell)},$ (32a)
$\boldsymbol{\Sigma}_{j\mid j-1}$ $=\sum_{\ell=0}^{2n}W^{(\ell)}_{c}\left(\mathcal{X}_{j\mid j-1}^{(\ell)}-\hat{\boldsymbol{x}}_{j\mid j-1}\right)\left(\mathcal{X}_{j\mid j-1}^{(\ell)}-\hat{\boldsymbol{x}}_{j\mid j-1}\right)^{T}+\boldsymbol{Q}_{j},$ (32b)
where $\boldsymbol{Q}_{j}$ is the process noise given by expression (28), while $W^{(\ell)}_{m}$ and $W^{(\ell)}_{c}$ are weights given by
$W^{(0)}_{m}$ $=\dfrac{\lambda}{n+\lambda},$ (33)
$W^{(0)}_{c}$ $=\dfrac{\lambda}{n+\lambda+(1-\alpha^{2}+\beta)},$
$W^{{(\ell)}}_{m}$ $=\dfrac{1}{2(n+\lambda)},\quad\ell=1,\ldots,2n,$
$W^{(\ell)}_{c}$ $=\dfrac{1}{2(n+\lambda)},\quad\ell=1,\ldots,2n,$
and $\beta$ is an additional parameter to incorporate prior knowledge of $\hat{\boldsymbol{x}}$; $\beta=2$ is optimal assuming the distribution of $\hat{\boldsymbol{x}}$ is Gaussian.